OpenStudy (anonymous):

What is the quotient (2x^2+10x+12)/(x+3)

OpenStudy (anonymous):

\[\frac{ 2\left( x ^{2}+5x+6 \right) }{\left( x+3 \right) }= \frac{ 2\left\{ x ^{2}+3x+2x+6 \right\} }{ \left( x+3 \right) }\] factorise and cancel.